ANALYSIS OF HIGH RESOLUTION INFRARED SPECTRA OF 1,1-DICHLOROETHYLENE IN THE 500− 1000 cm^{−1} RANGE

\underline{REBECCA A. PEEBLES}, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920; DANIEL A. OBENCHAIN, Department of Chemistry, Wesleyan University, 52 Lawn Avenue, Middletown, CT 06459-0180.

The far infrared beamline of the Canadian Light Source synchrotron facility has been used to record three rotationally resolved vibrational bands of 1,1-dichloroethylene in the 500 − 1000 cm^{−1} range, at 0.00096 cm^{−1} resolution. These correspond, for the H_{2}C=C^{35}Cl_{2} isotopologue, to an a-type band (CCl_{2} antisymmetric stretch) at 796.0 cm^{−1}, a b-type band (CCl_{2} symmetric stretch) at 603.0 cm^{−1}, and a c-type band (CH_{2} wag) at 868.6 cm^{−1}. Anharmonic frequency calculations at the MP2/6-311++G(2d,2p) level, combined with rotational and centrifugal distortion constants from a millimeter wave study of the ground state^{a}, were an invaluable aid in facilitating the spectroscopic assignment for this asymmetric top (κ = −0.58). Analysis of the 796 cm^{−1} band is nearly complete, giving well determined excited state rotational and centrifugal distortion constants. Results of this analysis and progress with analysis of the other two bands will be presented.

aZ. Kisiel, L. Pszczółkowski, Z. Naturforsch, \underline{50a}, (1995), 347-351.